OpenStudy (anonymous):

vertices of PQR are P(5,-2,-1),Q(1,2,-3) and R(13,6,3). Bisector of angle P meets QR at S. Find coordinates of S?

OpenStudy (anonymous):

angular bisector divides the opposite sides internally in the ratio of the side containing the angle

OpenStudy (anonymous):

so first find the length of PQ and PR and then divide QR in the rato of PQ :PR

OpenStudy (anonymous):

PQ= 6 PR = 12

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

so find the co-ordinates of a point S which divides QR in the ratio of 6:12 i.e 1:2 internally

OpenStudy (anonymous):

x=15/2

OpenStudy (anonymous):

y = 20/3 and z =-1 coordinates of S =15/2,20/3,-1

OpenStudy (anonymous):

@matricked am i right

OpenStudy (anonymous):

i don't think its the correct cordinate u obtained

OpenStudy (anonymous):

Q(1,2,-3) and R(13,6,3) check if it is (2+13)/3 , (4+6)/3 ,(-6+3)/3--> (15/3),(10/3),-1

OpenStudy (anonymous):

oh yes, i thought its 16 , but it was 6

OpenStudy (anonymous):

so u got it ..

OpenStudy (anonymous):

yes..and of course thank u so much @matricked

OpenStudy (anonymous):

welcome dear